# Constraint-Generation Policy Optimization (CGPO): Nonlinear Programming for Policy Optimization in Mixed Discrete-Continuous MDPs

### v1

@@ Abstract
We propose Constraint-Generation Policy Optimization (CGPO) for optimizing policy parameters within compact and interpretable policy classes for mixed discrete-continuous Markov Decision Processes (DC-MDPs). CGPO is not only able to provide bounded policy error guarantees over an infinite range of initial states for many DC-MDPs with expressive nonlinear dynamics, but it can also provably derive optimal policies in cases where it terminates with zero error. Furthermore, CGPO can generate worst-case state trajectories to diagnose policy deficiencies and provide counterfactual explanations of optimal actions. To achieve such results, CGPO proposes a bi-level mixed-integer nonlinear optimization framework for optimizing policies within defined expressivity classes (i.e. piecewise (non)-linear) and reduces it to an optimal constraint generation methodology that adversarially generates worst-case state trajectories. Furthermore, leveraging modern nonlinear optimizers, CGPO can obtain solutions with bounded optimality gap guarantees. We handle stochastic transitions through explicit marginalization (where applicable) or chance-constraints, providing high-probability policy performance guarantees. We also present a road-map for understanding the computational complexities associated with different expressivity classes of policy, reward, and transition dynamics. We experimentally demonstrate the applicability of CGPO in diverse domains, including inventory control, management of a system of water reservoirs, and physics control. In summary, we provide a solution for deriving structured, compact, and explainable policies with bounded performance guarantees, enabling worst-case scenario generation and counterfactual policy diagnostics.
Introduction
@@
An important aim of sequential decision optimization of challenging Discrete-Continuous Markov Decision Processes (DC-MDPs) in the artificial intelligence, operations research and control domains is to derive policies that achieve optimal control. A desirable property of such policies is compactness of representation, which provides efficient execution on resource-constrained systems such as mobile devices (Wang et al. 2022), as well as the potential for introspection and explanation (Topin et al. 2021). Moreover, while the derived policy is expected to perform well in expectation, in many applications it is desirable to obtain bounds on maximum policy error and and the scenarios that induce worst-case policy performance (Corso et al. 2021).
Existing approaches in the planning literature to inform the discovery of good plans or policies often leverage model structure, for instance by accelerating search (Keller and Eyerich 2012; Świechowski et al. 2023) or computing gradients (Wu, Say, and Sanner 2017; Bueno et al. 2019; Low, Kumar, and Sanner 2022), but do not generally guarantee optimality nor provide gap estimates for the performance of the plans or policies. On the other hand, existing methodologies for finding compact, optimal policies are often highly specialized to work in specific domains, such as linear–quadratic controllers with Gaussian noise (Åström 2012), ambulance dispatching (Albert 2022), and the “$(s,S)$” threshold policies in Scarf et al. (1960)’s formulation of the inventory management problem, or they provide only limited control over the degree of expressivity of the policy classes that can be optimized (Poupart and Boutilier 2003; Raghavan et al. 2017; Say et al. 2017; Topin et al. 2021; Vos and Verwer 2023). Meanwhile, symbolic dynamic programming (SDP) approaches can be used to derive symbolic representations of optimal policies in highly restricted subsets of DC-MDPs (Boutilier, Reiter, and Price 2001; Feng and Hansen 2002; Sanner, Delgado, and de Barros 2011; Raghavan et al. 2012; Zamani, Sanner, and Fang 2012), but do not often produce compact nor interpretable policies when applicable. To complicate matters, in more complex problems with nonlinear dynamics, a compact optimal policy may not exist at all, and even the function class containing the optimal policy may be unknown. In such settings, the ability to analyze the performance gap and interpret a compact policy is particularly beneficial for deciding upon the right expressivity class for policies.
@@
Figure 1: Reservoir control is used as an illustrative example. Left: an overview of CGPO, which consists of a domain description and policy representation compiled to a bilevel mixed-integer program (MIP), in which the inner problem computes the worst-case trajectories for the current policy while the outer problem updates the policy via constraint-generation. The result is a worst-case scenario, a concrete policy within the expressivity class, and a gap on its performance. Right: three optimal (i.e. zero-gap) policies produced upon termination across several piecewise policy classes. Crucially, our framework provides the ability to derive highly compact (e.g. memory and time-efficient to execute), intuitive and nonlinear policies, with strong bound guarantees on policy performance.
In this paper, we aim to provide a solution approach to optimize and provide strong performance bound guarantees on structured and compact DC-MDP policies under various expressivity classes of policy, reward, and (nonlinear) transition dynamics. Our specific contributions are as follows:
@@
    We propose a novel nonlinear bi-level optimization framework called Constraint-Generation Policy Optimization (CGPO), that admits a clever reduction to an iterative constraint-generation algorithm that can be readily implemented using standard MIP solvers (cf. Figure 1). State-of-the-art nonlinear solvers often leverage spacial branch-and-bound techniques, which can provide not only optimal solutions for large mixed integer linear programs (MILP), but also bounded optimality guarantees for many cases of mixed integer nonlinear programs (MINLP) (Castro 2015). We efficiently handle stochastic DC-MDPs using marginalization, or chance-constraints (Farina, Giulioni, and Scattolini 2016) to provide probabilistic guarantees on policy performance.
@@
    If the constraint generation algorithm terminates, we provide a guarantee that we have found the optimal policy within the given policy expressivity class (Theorem 1). Further, even when constraint generation does not terminate, our algorithm provides a tight optimality bound on the performance of the computed policy at each iteration and the scenario where the policy performs worst (and as a corollary, for any externally provided policy).
@@
    We provide a road map to characterize the optimization problems in (1) above – and their associated expressivity classes – for different expressivity classes of policies and state dynamics, ranging from MILP, to polynomial programming, to nonlinear programs (cf. Table 1). This information is beneficial for reasoning about which optimization techniques are most effective for different combinations of DC-MDP and policy class expressivity.
@@
    Finally, we provide a variety of experiments demonstrating the range of rich applications of CGPO under linear and nonlinear dynamics and various policy classes. Critically, we derive bounded optimal solutions for a range of problems, including linear inventory management, reservoir control, and a highly nonlinear VTOL control problem. Further, since our policy classes are compact by design, we can also directly inspect and analyze these policies (cf. Figure 2). To facilitate policy interpretation and diagnostics, we can compute the state and exogenous noise trajectory scenario that attains the worst-case error bounds for a policy (cf. Figure 5) as well as a counterfactual explanation of what action should have been taken in comparison to what the policy prescribes.
Preliminaries
Function Classes
@@
We first define some important classes of functions commonly referred to throughout the paper. Let $\mathcal{X}$ and $\mathcal{Y}$ be arbitrary sets. A function $f:\mathcal{X}\to\mathcal{Y}$ is called piecewise (PW) if there exist functions $f_{1},\dots f_{K},f_{K+1}:\mathcal{X}\to\mathcal{Y}$ and disjoint Boolean predicates $\mathcal{P}_{1}\dots\mathcal{P}_{K}:\mathcal{X}\to\{0,1\}$, such that for any $x\in\mathcal{X}$, $\forall i,j\!\neq\!i\,\,\mathcal{P}_{i}(x)\!=\!1\,\Rightarrow\,\mathcal{P}_{j}(x)\!=\!0$ and
$f(x)=\begin{cases}f_{1}(x),&\mbox{ if }\mathcal{P}_{1}(x)=1,\\ &\vdots\\ f_{K}(x)&\mbox{ if }\mathcal{P}_{K}(x)=1,\\ f_{K+1}(x)&\mbox{ otherwise}.\end{cases}$ (1)
@@
Both predicates $\mathcal{P}_{i}$ and cases ${f}_{i}$ can be either linear or nonlinear. In this paper, we consider the following classes:
  • constant (C), i.e. $f_{i}(\mathbf{x})=C$
  • discrete (D), same as restricted values of C
@@
  • simple (axis-aligned) linear functions (S), i.e. $f_{i}(\mathbf{x})=b_{i}+w_{i}\times x_{j}$, where $j\in\{1,2\dots n\}$ and $b_{i},w_{i}\in\mathbb{R}$
  • linear (L), i.e. $f_{i}(\mathbf{x})=b_{i}+\mathbf{w}_{i}^{T}\mathbf{x}$
  • bilinear (B), i.e. $f_{i}(\mathbf{x})=x_{1}\times x_{2}+x_{2}\times x_{3}$
  • quadratic (Q), i.e. $f_{i}(\mathbf{x})=b_{i}+\mathbf{w}_{i}^{T}\mathbf{x}+\mathbf{x}^{T}M_{i}\mathbf{x}$, where $M_{i}\in\mathbb{R}^{n\times n}$
  • polynomial (P) of order $m$, i.e. $f_{i}(\mathbf{x})=\sum_{j_{1}=1}^{m}\dots\sum_{j_{n}=1}^{m}w_{i,j_{1},\dots j_{n}}\prod_{k=1}^{n}x_{k}^{j_{k}}$
  • general nonlinear (N), i.e. $f(\mathbf{x})=e^{b_{i}+\mathbf{w}_{i}^{T}\mathbf{x}}$.
@@
In a similar way, we can also consider analogous functions over integer (I) domains $\mathcal{X}$ or $\mathcal{Y}$, or mixed discrete-continuous (M) domains. $\mathcal{P}_{i}$ can be characterized in a similar way based on whether the constraint, or set of constraints, defining it are constant, linear, bilinear, quadratic, etc.
@@
We also introduce a convenient notation to describe general PW functions by concatenating the expressivity classes of their constraints $\mathcal{P}_{i}$ and case functions $f_{i}$. For example, PWS-C describes a piecewise function on $\mathbb{R}^{n}$ with univariate axis-aligned constraints and constant values, while PWS-L describes a similar function but with linear values, i.e.:
$f(\mathbf{x})=\begin{cases}b_{1}+\mathbf{w}_{1}^{T}\mathbf{x},&\mbox{ if }l_{1}\leq c_{1}+v_{1}x_{i_{1}}\leq u_{1},\\ &\vdots\\ b_{K}+\mathbf{w}_{K}^{T}\mathbf{x}&\mbox{ if }l_{K}\leq c_{K}+v_{K}x_{i_{K}}\leq u_{K},\\ b_{K+1}+\mathbf{w}_{K+1}^{T}\mathbf{x}&\mbox{ otherwise}.\end{cases}$
We also append the number of cases $K$ to the notation, i.e. PWS1-L describes a piecewise linear function as above with $K=1$, and we assume that $K$ is finite in our derivation. More generally, for functions in PWL, PWP and PWN, constraints could also be conjunctions of other simpler constraints, i.e. $\mathcal{P}_{i}(\mathbf{x})=\bigwedge\limits_{j=1}^{c}\left\{l_{ij}\leq b_{ij}+\mathbf{w}_{ij}^{T}\mathbf{x}\leq u_{ij}\right\}$.
Mathematical Programming
@@
Given a function $f:\mathcal{X}\to\mathcal{Y}$ and Boolean predicate $\mathcal{P}:\mathcal{X}\to\{0,1\}$, we define the general mathematical program:
[…]
@@
Important standard expressivity classes of mixed-integer program (MIP) optimization problems, i.e. those with discrete and continuous decision variables, include:
  • mixed-integer linear (MILP), i.e. $f$, $\mathcal{P}$ are both L
  • mixed-integer bilinearly constrained (MIBCP), i.e. $f$ is L and $\mathcal{P}$ is B
  • mixed-integer quadratic (MIQP), i.e. $f$ is Q and $\mathcal{P}$ is L
  • quadratically constrained quadratic (QCQP), i.e. $f$, $\mathcal{P}$ are both Q
  • polynomial (PP), i.e. $f$, $\mathcal{P}$ are both P
  • mixed-integer nonlinear (MINLP), i.e. $f$, $\mathcal{P}$ are both N.
@@
Branch-and-bound (Morrison et al. 2016) is a powerful search technique for finding optimal solutions to mathematical programs, and is a key component of modern state-of-the-art nonlinear optimization software. One advantage of this class of algorithms is the ability to maintain upper and lower bounds on the minimal objective value of any linear or nonlinear MIP (and thus the optimal solution in any branch of the search tree), whose difference is called the optimality gap; when this gap is zero, an optimal solution has been found. Moreover, some packages such as Gurobi (Gurobi Optimization, LLC 2023), which we use to perform the necessary compilations in CGPO in our experiments, also support a variety of nonlinear mathematical operations via piecewise-linear approximation (Castro 2015).
Discrete-Continuous Markov Decision Processes
@@
A Discrete-Continuous Markov decision process (DC-MDP) is a tuple $\langle\mathcal{S},\mathcal{A},P,r\rangle$, where $\mathcal{S}$ is a set of states, $\mathcal{A}$ is a set of actions or controls. $\mathcal{S}$ and $\mathcal{A}$ may be discrete, continuous, or mixed. $P(\mathbf{s},\mathbf{a},\mathbf{s}^{\prime})$ is the probability of transitioning to state $\mathbf{s}^{\prime}$ immediately upon choosing action $\mathbf{a}$ in state $\mathbf{s}$, and $r(\mathbf{s},\mathbf{a},\mathbf{s}^{\prime})$ is the corresponding reward received. We assume that $r$ is uniformly bounded, i.e. there is a $B<\infty$ such that $r(\mathbf{s},\mathbf{a},\mathbf{s}^{\prime})\leq B$ holds for all $\mathbf{s},\mathbf{a},\mathbf{s}^{\prime}$.
@@
A policy $\pi=[\pi_{1},\dots\pi_{T}]$ is a sequence of mappings $\pi_{t}:\mathcal{S}\to\mathcal{A}$, whose value is defined as
@@
Dynamic programming approaches such as value or policy iteration can compute an optimal horizon-dependent policy $\pi_{H}^{*}$ (Puterman 2014), but do not directly apply to DC-MDPs with infinite or continuous state or action spaces.
Our goal is to compute an optimal stationary reactive policy $\pi^{*}$ (Bueno et al. 2019) that minimizes the error in value (i.e. regret) relative to $\pi_{H}^{*}$ over all initial states of interest $\mathcal{S}_{1}\subseteq\mathcal{S}$, and all stationary policies $\Pi$, i.e.
@@
In practical applications, the policy class is typically restricted to some class of function approximations $\tilde{\Pi}\subset\Pi$. A variety of planning approaches can compute the optimal policies $\tilde{\pi}^{*}\in\tilde{\Pi}$ for this problem, including straight-line planning that scales well in practice but does not learn policies (Wu, Say, and Sanner 2017; Raghavan et al. 2017), and deep reactive policy (Bueno et al. 2019; Low, Kumar, and Sanner 2022), MCTS (Świechowski et al. 2023) and similar approaches that learn neural network policies but cannot provide concrete bounds on performance, nor compactness or ease of interpretation of the inferred policy map.
Methodology
Our main contribution is a novel iterative constraint generation algorithm for policy optimization with structured (i.e. piecewise) policies. We begin with a simple worked example to illustrate the intuition behind our approach. Next, we present the constraint generation approach for general deterministic DC-MDPs, and then describe the chance-constraints modification that works with high probability for stochastic DC-MDPs.
Worked Example
In order to shed some intuition on our constraint-generation approach in the general MDP setting, we illustrate our approach in the context of a one-dimensional “navigation” type problem.
Domain Description
The Outer Problem
For ease of illustration, we focus our attention on linear policies of the form $a_{t}=b+ws_{t}$, where $(w,b)$ are real-valued decision variables to be optimized. The goal is to find the optimal policy parameters:
$\min_{(w,b)\in\mathbb{R}^{2},\,\varepsilon\in[0,\infty)}\varepsilon,\quad\textrm{s.t.}\quad\varepsilon\geq V(a_{1},s_{1})-V((w,b),s_{1}),$
where constraints are enumerated by all possible $(s_{1},a_{1})$ pairs. Since the number of constraints is infinite for continuous states or actions, it suggests a constraint generation approach, where a large set of diverse $(s_{1},a_{1})$ pairs are used to “build up” the constraint set one at a time, decoupling the overall problem into a sequence of solvable MIPs. However, a key question arises: which state-action pairs $(s_{1},a_{1})$ should be considered for inclusion into the outer problem? While a number of different reasonable choices (diversity, coverage) exist, a logical choice is to select the state action pairs in relation to which the policy $\pi_{\mathbf{w}}$ performs worst.
Solving the Outer Problem
The outer problem thus maintains a finite subset of constraints from the infinitely-constrained problem. We begin with an arbitrary $(s,a)$-pair, which forms the first constraint of the outer problem. For illustration, we select $(s_{1},a_{1})=(0,0)$, which defines the first constraint of the outer problem $\varepsilon\geq V(0,0)-V((w,b),0)=-10+|b-10|$. The outer problem becomes:
$\min_{w,b\in\mathbb{R},\,\varepsilon\in[0,\infty)}\,\varepsilon,\quad\mathrm{s.t.}\quad\varepsilon\geq-10+|b-10|.$
Solving the Inner Sub-Problem
To accomplish this, we need to add a new scenario $(s_{1},a_{1})$ in relation to which the policy $(w,b)=(0,10)$ performs worst. Thus, we solve:
$\max_{s_{1}\in[0,5],a_{1}\in\mathbb{R}}[V(a_{1},s_{1})-V((w,b),s_{1})]$
$=\max_{s_{1}\in[0,5],a_{1}\in\mathbb{R}}[-|s_{1}+a_{1}-10|+s_{1}],$
Solving the Outer Problem Again
We add the constraint corresponding to this scenario to the outer problem:
$\min_{(w,b)\in\mathbb{R}^{2},\,\varepsilon\in[0,\infty)}\quad\varepsilon\qquad\textrm{s.t.}\quad\varepsilon$ $\geq-10+|b-10|$
$\varepsilon$ $\geq|b+5w-5|,$
whose new solution can be found as $b^{*}=10,w^{*}=-1$ with optimal $\varepsilon=0$. Finally, the worst-case scenario for this new policy can be found from the inner sub-problem, whose optimal solution (among many) is $(s_{1}^{*},a_{1}^{*})=(0,10)$ with value zero. Therefore, this worst-case corresponds to a non-binding constraint in the outer problem, and we terminate with the optimal linear policy $\pi^{*}(s_{1})=10-s_{1}$. $\square$
Constraint Generation for Deterministic DC-MDPs

### v2

# Bounded-Error Policy Optimization for Mixed Discrete-Continuous MDPs via Constraint Generation in Nonlinear Programming

@@ Abstract
We propose the Constraint-Generation Policy Optimization (CGPO) framework to optimize policy parameters within compact and interpretable policy classes for mixed discrete-continuous Markov Decision Processes (DC-MDP). CGPO can not only provide bounded policy error guarantees over an infinite range of initial states for many DC-MDPs with expressive nonlinear dynamics, but it can also provably derive optimal policies in cases where it terminates with zero error. Furthermore, CGPO can generate worst-case state trajectories to diagnose policy deficiencies and provide counterfactual explanations of optimal actions. To achieve such results, CGPO proposes a bilevel mixed-integer nonlinear optimization framework for optimizing policies in defined expressivity classes (e.g. piecewise linear) and reduces it to an optimal constraint generation methodology that adversarially generates worst-case state trajectories. Furthermore, leveraging modern nonlinear optimizers, CGPO can obtain solutions with bounded optimality gap guarantees. We handle stochastic transitions through chance constraints, providing high-probability performance guarantees. We also present a roadmap for understanding the computational complexities of different expressivity classes of policy, reward, and transition dynamics. We experimentally demonstrate the applicability of CGPO across various domains, including inventory control, management of a water reservoir system, and physics control. In summary, CGPO provides structured, compact and explainable policies with bounded performance guarantees, enabling worst-case scenario generation and counterfactual policy diagnostics.
Keywords Planning  $\cdot$ Control  $\cdot$ Mixed Discrete-Continuous MDP  $\cdot$ Constraint Generation  $\cdot$ Chance Constraints  $\cdot$ Piecewise-Linear Policy  $\cdot$ Sequential Decision Optimization  $\cdot$ Policy Optimization
1 Introduction
@@
An important aim of sequential decision optimization of challenging Discrete-Continuous Markov Decision Processes (DC-MDP) in the artificial intelligence, operations research, and control domains is to derive policies that achieve optimal control. A desirable property of such policies is compactness of representation, which provides efficient execution on resource-constrained systems such as mobile devices (Wang et al., 2022), as well as the potential for introspection and explanation (Topin et al., 2021). Moreover, while the derived policy is expected to perform well in expectation, in many applications it is desirable to obtain bounds on maximum policy error and the scenarios that induce worst-case policy performance (Corso et al., 2021).
Popular policy optimization approaches used in model-free reinforcement learning (Schulman et al., 2017) do not provide bounded error guarantees on policy performance, even if they use compact policy classes such as decision trees (Topin et al., 2021). Model-based extensions that leverage gradient-based policy optimization (Bueno et al., 2019) similarly do not provide bounded error guarantees due to the nonconvexity of the problem.
In contrast, there exists a rich tradition of work leveraging mixed integer programming (MIP) for bounded-error policy optimization (Ahmed et al., 2017; Dolgov and Durfee, 2005; Vos and Verwer, 2023), but these methods only apply to discrete state and action MDPs. More importantly, these previous policy optimization formulations cannot be directly extended to continuous state or action MDPs due to the infinite number of constraints arising from the infinite state and action space in these formulations. While some historical work has attempted to circumvent large or infinite constraint sets (from the continuous setting) via constraint generation (Farias and Van Roy, 2006; Hauskrecht, 2006; Schuurmans and Patrascu, 2001), none of these works optimize for policies. Finally, while there exist bespoke bounded error policy solutions for niche MDP classes with continuous states or actions such as linear–quadratic controllers (Åström, 2012), ambulance dispatching (Albert, 2022), and the celebrated “$(s,S)$” threshold policies for (Scarf et al., 1960)’s inventory management, these solution methods do not readily generalize beyond the specialized domains for which they are designed.
It remains an open question as to how to derive bounded error policies for the infinite state and action, discrete and continuous MDP (DC-MDP) setting. We address this open question through a novel bilevel MIP formulation. Unfortunately, unlike MIP solutions for discrete MDPs, DC-MDPs with continuous states and/or actions cannot be solved directly since the resulting MIP (a) requires a challenging bilevel MIP formulation, (b) has an uncountably infinite number of constraints, and (c) will often be nonlinear. Fortuitously, we show how a constraint generation form of this policy optimization – termed CGPO – is able to address (a), (b) and (c) to optimize policies with bounded error guarantees in many cases.
@@
Figure 1: Reservoir control is used as an illustrative example. Left: an overview of CGPO, which consists of a domain description and policy representation compiled to a bilevel mixed-integer program (MIP), in which the inner problem computes the worst-case trajectories for the current policy while the outer problem updates the policy via constraint-generation. The result is a worst-case scenario for the policy (facilitating policy failure analysis), a concrete policy within the expressivity class (for direct policy inspection), and a gap on its performance (error bound). Right: three optimal (i.e. zero-gap) policies produced upon termination across several piecewise policy classes. Crucially, our framework provides the ability to derive highly compact (e.g. memory and time-efficient to execute), intuitive and nonlinear policies, with strong bound guarantees on policy performance.
In summary, our aim is to provide a solution approach to optimize and provide strong performance bound guarantees on structured and compact DC-MDP policies under various expressivity classes of policy, reward, and nonlinear transition dynamics. Our specific contributions are:
@@
    We propose a novel bilevel optimization framework for solving nonlinear DC-MDPs called Constraint-Generation Policy Optimization (CGPO), that admits a clever reduction to an iterative constraint-generation algorithm that can be readily implemented using standard MIP solvers (Fig. 1). State-of-the-art MIP solvers often leverage spatial branch-and-bound techniques, which can provide not only optimal solutions for large mixed integer linear programs (MILP), but also bounded optimality guarantees for mixed integer nonlinear programs (MINLP) (Castro, 2015). Chance constraints (Farina et al., 2016) are used for probabilistic guarantees.
@@
    If the constraint generation algorithm terminates, we guarantee that we have found the optimal policy within the given policy expressivity class (Theorem 1). Further, even when constraint generation does not terminate, our algorithm provides a tight optimality bound on the performance of the computed policy at each iteration and the scenario where the policy performs worst (and as a corollary, for any externally provided policy).
@@
    We provide a road map to characterize the optimization problems in (1) above – and their associated expressivity classes – for different expressivity classes of policies and state dynamics, ranging from MILP, to polynomial programming, to nonlinear programs (Table 1). This information is beneficial for reasoning about which optimization techniques are most effective for different combinations of DC-MDP and policy class expressivity.
@@
    Finally, we provide a variety of experiments demonstrating the range of rich applications of CGPO under linear and nonlinear dynamics and various policy classes. Critically, we derive bounded optimal solutions for a range of problems, including linear inventory management, reservoir control, and a highly nonlinear VTOL control problem. Furthermore, since our policy classes are compact by design, we can also directly inspect and analyze these policies (Fig. 4). To facilitate policy interpretation and diagnostics, we can compute the state and exogenous noise trajectory scenario that attains the worst-case error bounds for a policy (Fig. 7) as well as a counterfactual explanation of what action should have been taken in comparison to what the policy prescribes.
2 Preliminaries
2.1 Function Classes
@@
We first define some important classes of functions commonly referred to throughout the paper. Let $\mathcal{X}$ and $\mathcal{Y}$ be arbitrary sets. A function $f:\mathcal{X}\to\mathcal{Y}$ is called piecewise (PW) if there exist functions $f_{1},\dots f_{K},f_{K+1}:\mathcal{X}\to\mathcal{Y}$ and disjoint Boolean case conditions $\mathcal{P}_{1},\dots\mathcal{P}_{K}:\mathcal{X}\to\{0,1\}$, such that for any $x\in\mathcal{X}$, $\forall i,j\!\neq\!i$ we have $\mathcal{P}_{i}(x)\!=\!1\,\Rightarrow\,\mathcal{P}_{j}(x)\!=\!0$, and for case functions $f(x)=f_{i}(x)$ if $\mathcal{P}_{i}(x)=1$ and $f(x)=f_{K+1}(x)$ if $\mathcal{P}_{i}(x)=0,\,\forall i$.
@@
Both case conditions $\mathcal{P}_{i}$ and case functions ${f}_{i}$ can be linear or nonlinear. In this paper, we consider the following classes:
C
constant, i.e. $f_{i}(\mathbf{x})=C$
D
discrete, same as restricted values of C
S
@@
simple (axis-aligned) linear functions, i.e. $f_{i}(\mathbf{x})=b_{i}+w_{i}\times x_{j}$, where $j\in\{1,2\dots n\}$ and $b_{i},w_{i}\in\mathbb{R}$
L
linear, i.e. $f_{i}(\mathbf{x})=b_{i}+\mathbf{w}_{i}^{T}\mathbf{x}$
B
bilinear, i.e. $f_{i}(\mathbf{x})=x_{1}\times x_{2}+x_{2}\times x_{3}$
Q
quadratic, i.e. $f_{i}(\mathbf{x})=b_{i}+\mathbf{w}_{i}^{T}\mathbf{x}+\mathbf{x}^{T}M_{i}\mathbf{x}$, where $M_{i}\in\mathbb{R}^{n\times n}$
P
polynomial of order $m$, i.e. $f_{i}(\mathbf{x})=\sum_{j_{1}=1}^{m}\dots\sum_{j_{n}=1}^{m}w_{i,j_{1},\dots j_{n}}\prod_{k=1}^{n}x_{k}^{j_{k}}$
N
general nonlinear, i.e. $f(\mathbf{x})=e^{b_{i}+\mathbf{w}_{i}^{T}\mathbf{x}}$.
@@
Analogously, we consider analogous functions over integer (I) domains $\mathcal{X}$ or $\mathcal{Y}$, or mixed discrete-continuous (M) domains. $\mathcal{P}_{i}$ can be characterized similarly based on whether the constraint set defining it is constant, linear, bilinear, etc.
@@
We also introduce a convenient notation to describe general PW functions by concatenating the expressivity classes of their constraints $\mathcal{P}_{i}$ and case functions $f_{i}$. For example, PWS-C describes a piecewise function on $\mathbb{R}^{n}$ with simple (axis-aligned) linear constraints and constant values, while PWS-L describes a function with similar constraints but linear values. We can also append the number of cases $K$ to the notation, i.e. PWS1-L describes a piecewise linear function with $K=1$ case conditions. More generally, for PWL, PWP or PWN, constraints could also be logical conjunctions of other simpler constraints.
2.2 Mathematical Programming
@@
Given a function $f:\mathcal{X}\to\mathcal{Y}$ and Boolean case condition $\mathcal{P}:\mathcal{X}\to\{0,1\}$, the mathematical program is defined as:
$\min_{x\in\mathcal{X}}\quad f(x)\qquad\textrm{s.t.}\quad\mathcal{P}(x)=1.$
@@
Important expressivity classes of mixed-integer program (MIP) optimization problems, i.e. those with discrete and continuous decision variables, include:
MILP
mixed-integer linear: $f$ and $\mathcal{P}$ both in L
MIBCP
mixed-integer bilinearly constrained: $f$ is in L and $\mathcal{P}$ is in B
MIQP
mixed-integer quadratic: $f$ is in Q and $\mathcal{P}$ is in L
QCQP
quadratically constrained quadratic: $f$ and $\mathcal{P}$ both in Q
PP
polynomial: $f$ and $\mathcal{P}$ both in P
MINLP
mixed-integer nonlinear: $f$ and $\mathcal{P}$ both in N.
@@
Branch-and-bound (Morrison et al., 2016) solvers are commonly used to maintain upper and lower bounds on the minimal objective value of any linear or nonlinear MIP, whose difference is the so-called optimality gap; when this gap is zero, an optimal solution has been found. Moreover, some packages such as Gurobi, which we use to perform the necessary compilations in CGPO in our experiments, also support a variety of nonlinear mathematical operations via piecewise-linear approximation (Castro, 2015).
2.3 Discrete-Continuous Markov Decision Processes
@@
A Discrete-Continuous Markov decision process (DC-MDP) is a tuple $\langle\mathcal{S},\mathcal{A},P,r\rangle$, where $\mathcal{S}$ is a set of states, $\mathcal{A}$ is a set of actions or controls. $\mathcal{S}$ and $\mathcal{A}$ may be discrete, continuous, or mixed. $P(\mathbf{s},\mathbf{a},\mathbf{s}^{\prime})$ is the probability of transitioning to state $\mathbf{s}^{\prime}$ immediately upon choosing action $\mathbf{a}$ in state $\mathbf{s}$, and $r(\mathbf{s},\mathbf{a},\mathbf{s}^{\prime})$ is the corresponding reward received. We assume that $r$ is uniformly bounded, i.e. there exists $B<\infty$ such that $|r(\mathbf{s},\mathbf{a},\mathbf{s}^{\prime})|\leq B$ holds for all $\mathbf{s},\mathbf{a},\mathbf{s}^{\prime}$.
@@
A (closed-loop) policy $\pi=[\pi_{1},\dots\pi_{T}]$ is a sequence of mappings $\pi_{t}:\mathcal{S}\to\mathcal{A}$, whose value is defined as
@@
Dynamic programming approaches such as value and policy iteration can compute an optimal horizon-dependent policy $\pi_{H}^{*}$ (Puterman, 2014), but do not directly apply to DC-MDPs with infinite or continuous state or action spaces.
Our goal is to compute an optimal stationary policy $\pi^{*}$ that minimizes the error in value relative to $\pi_{H}^{*}$ over all initial states of interest $\mathcal{S}_{1}\subseteq\mathcal{S}$, and all stationary policies $\Pi$, i.e.
@@
In practical applications, the policy class is often restricted to function approximations $\tilde{\Pi}\subset\Pi$. A variety of planning approaches can compute plans $\tilde{\pi}^{*}\in\tilde{\Pi}$ for this problem, including straight-line planning that scales well in practice but does not learn policies (Raghavan et al., 2017; Wu et al., 2017). Reactive policy optimization (Bueno et al., 2019; Low et al., 2022) and variants of MCTS (Świechowski et al., 2023) that learn neural network policies cannot provide concrete error bounds on policy performance nor do they facilitate worst-case analysis, or ease of interpretation of learned compact policy classes as we provide.
3 Methodology
We begin with a derivation of CGPO for general DC-MDPs in the deterministic setting. We then derive CGPO in the more general stochastic setting using chance constrained optimization. We end with a discussion of problem class complexity and convergence. A worked example illustrating the overall execution of CGPO is provided in Appendix 6.1.
3.1 Constraint Generation for Deterministic DC-MDPs